OpenStudy (anonymous):

Need Help: Qd = 17000 –6P Qs = 900 + 8P Find out the price elasticity of demand and price elasticity of supply of Sports goods when the industry is in equilibrium and interpret the results.

OpenStudy (anonymous):

in order to be able to calculate the elasticity we need to know what is the price and quantity. lets calculate ;; QD=QS 17000-6P=900+8P 17000-900=8P+6P 16100=14P 16100/14=14P/14 1150=P 10100=Q E=Q2-Q1/Q2/P2-P1/P2=10100-10100/10100 /1150-1150/1150 =0 THE PRICE ELASTICITY OF DEMAND AND SUPPLY IN EQ IS PERFECTLY IENELASTICS ANY CHANGE IN EQUELIBRIUM TO PRICE OR QUANTITY WILL NOT HAVE ANY EFFECT TO PRICE VISE VERSA

OpenStudy (anonymous):

THanks